OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin 3θ cos θ − cos 3θ sin θ = sqrt(3)/2

OpenStudy (anonymous):

there is a problem on here that is almost the same so i will show you that one \[\sin θ \cos 3θ + \cos θ \sin 3θ = \sin (θ+3θ) = \sin 4θ\]